beryllium-11 decomposes into boron-11 with a half life of 13.8 seconds. How long will it take 240 g of beryllium-11 to decompose into 7.5g of beryllium-11?
based on the forumula provided in the text book C(t)=C(0)2e^-t/h
where C(0) = 240, C(t)=7.5 ,h=13.8
7.5/240=2e^-t/13.8
this was a s far as i got

Answer by stanbon(75887)

beryllium-11 decomposes into boron-11 with a half life of 13.8 seconds. How long will it take 240 g of beryllium-11 to decompose into 7.5g of beryllium-11?
based on the forumula provided in the text book C(t)=C(0)2e^-t/h
where C(0) = 240, C(t)=7.5 ,h=13.8
7.5/240=2e^-t/13.8
this was a s far as i got
------------------------------------
7.5/240=2e^(-t/13.8)
---
(7.5/480) = e^(-t/13.8)
Take the natural log of both sides to get:
-t/13.8 = -4.1589
t = 57.39 seconds
=====================
2nd attempt using a standard half-life formula:
A(t) = A(0)(1/2)^(t/h)
7.5 = 240*(1/2)^(t/13.8)
(7.5/240) = (1/2)^(t/13.8)
0.03125 = (1/2)^(t/13.8)
---
Take the log of both sides to get:
(t/13.8)*log(1/2) = log(0.03125)
---
(t/13.8) = -1.5052/-0.3010
t/13.8 = 5
t = 69 seconds
====================
Cheers,
Stan H.